what is the answer when we subtract 30 1/2 and 15 3/5
Answer by Edwin McCravy(20081)

 30 1/2 =  30 5/10 =  29 15/10
-15 3/5 = -15 6/10 = -15  6/10
                      14  9/10

In the first step, we got the LCD of 10.  Then we would have subtracted
if the larger fraction had been on the top, but since the smaller
fraction is on the top, we had to take another step to borrow 1 from 
the 30 and make it 29, then take that 1 that we just borrowed, think of it
as 10/10 and add 10/10 to the 5/10 fraction part making it 15/10,
so now the bigger fraction is on top and we can finally do the
subtracting.

Edwin